For the following, assume that all the given angles are in simplest form, so that if A is in QIV you may assume that 270° < A < 360°.
If sin B = − 1/5 with B in QIII, find the following.cot B/2

Answer by stanbon(75887)

If sin B = − 1/5 with B in QIII, find the following.
cot(B/2) = 1/tan(B/2) = (1 + sqrt(1+tan^2(B))/tan(B)
-----
Since sin(B) = -1/5, y = -1 and r = 5
Then x = sqrt(r^2-y^2) = sqrt(25-1) = sqrt(24)
Note: x is negative in QIII
So x = -sqrt(24)
============
Therefore tan(B) = y/x = -1/-sqrt(24) = 1/sqrt(24)
------------------------
Finally, cot(B/2) = (1 + sqrt(1+(1/24))/[1/sqrt(24)]
---
= (1+sqrt(25/24))*(sqrt(24))
---
= (1+5/(sqrt(24))(sqrt(24)
---
= sqrt(24)+5
======================
Cheers,
Stan H.